A car of mass 2500 kg and moving at 3 m/s, northeast collides with 3800 kg car moving northwest at 2 m/s. Find their velocity, assuming that the collision is perfectly inelastic.

1
Expert's answer
2022-01-26T17:46:12-0500
"Ox: m_1v_1=(m_1+m_2)u\sin\theta,\\
Oy: m_2v_2=(m_1+m_2)u\cos\theta,\\\space\\
\theta=44.6°,\\
u=1.69\text{ m/s}."